# Manifolds with kinks and the asymptotic behavior of the graph Laplacian operator with Gaussian kernel

### v2

@@ 3.3.2. Inward sector
Let us now introduce the notion of inward tangent sector for smooth manifolds with kinks. This is analogous to the inward tangent sector of Euclidean domains as discussed in Section 2.2. Our definition builds upon the classical characterization of the tangent space in terms of initial velocities of curves, which holds true for manifolds without/with boundary (see e.g. [Lee12, p. 68–70]). For manifolds with boundary, one can identify inward tangent vectors at boundary points by specifying the domain of the curves we choose. This is how we came up with the following natural definition.
@@ 3.4.4. Riemannian distance and volume measure
The Riemannian distance defined via a length-minimizing problem extends with no change to the context of manifolds with kinks. We recall the definition for completeness and refer to [Lee12, p.337–341], for instance, for more details.
@@ 3.5. Extension of Riemannian manifolds with kinks
For our purposes, we need to extend beyond the border any $\mathcal{C}^{2}$ Riemannian metric defined on a smooth manifold with kinks $M$. To this aim, we shall flow $M$ into its interior using a suitable semiflow. We adapt an argument for manifolds with corners that goes back to [DH73] at least, see also [Mic80, Section 2.7].
@@ 3.5.2. Riemannian submanifolds with kinks
Let us now provide a definition of submanifold adapted to the context of manifolds with kinks, inspired by [Mic80, p.19] who introduced submanifolds with corners.
@@
    Use Whitney’s theorem [Whi34] to extend each coordinate of $(\phi^{-1})^{*}g$ to form a $\mathcal{C}^{2}$ Riemannian metric $h$ on an open subset $V$ of $\mathbb{R}^{d}$ containing $\phi(x)$, such that $\phi^{-1}(V)\subset\tilde{U}$. Note that this extension may not be unique and depends on the choice of the border chart $(U,\phi)$.
@@ 4. Asymptotic behavior of the intrinsic Gaussian Operator
In this section, we prove Theorem 1.1.
@@ 4.1. Localisation
Recall that for a metric measure space $(M,\mathop{}\!\mathrm{d},\mu)$, the intrinsic $d$-dimensional Gaussian operator at time $t>0$ associated with a density $p$ is defined by
$L_{t}f(x):=\frac{1}{t^{d/2+1}}\int_{M}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(x,y)}{t}\right)(f(x)-f(y))p(x)\mathop{}\!\mathrm{d}\mu(x)$
for any $f\in\mathcal{C}(M)\cap L^{1}(M,p\mu)$ and $x\in M$. We are interested in the short-time behavior of this operator. In this regard, the next lemma implies that all the information about this behavior is contained in small balls. The proof is immediate so we skip it.
@@
Let $(M,\mathop{}\!\mathrm{d},\mu)$ be a metric measure space and $\eta\in(0,1/2)$. Then for any $t>0$,
$\left|\frac{1}{t^{d/2+1}}\int_{M\backslash B_{t^{\eta}}(x)}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(x,y)}{t}\right)(f(x)-f(y))p(y)\mathop{}\!\mathrm{d}\mathrm{\mu}(y)\right|\leq[|f(x)|\|p\|_{1}+\left\lVert fp\right\rVert_{1}]\,\frac{1}{t^{d/2+1}}e^{-t^{2\eta-1}}$
for any $f\in\mathcal{C}(M)\cap L^{1}(M,p\mu)$ and $x\in X$.
As a consequence, we get that
$L_{t}f(x)=\frac{1}{t^{d/2+1}}\int_{B_{t^{\eta}}(x)}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(x,y)}{t}\right)(f(x)-f(y))p(y)\mathop{}\!\mathrm{d}\mathrm{\mu}(y)+O(t^{-d/2-1}e^{-t^{2\eta-1}})$
as $t\downarrow 0$, hence Theorem 1.1 follows from studying the asymptotic expansion of the first term in the previous right-hand side. Let us give a name to this quantity:
(11) $\bar{L}_{t,\eta}f(x):=\frac{1}{t^{d/2+1}}\int_{B_{t^{\eta}}(x)}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(x,y)}{t}\right)(f(x)-f(y))p(y)\mathop{}\!\mathrm{d}\mathrm{\mu}(y)$
4.2. Change of coordinates
Let us put ourselves in the context of Theorem 1.1. We consider a smooth $d$-dimensional Riemannian manifold endowed with a $\mathcal{C}^{2}$ Riemannian metric, a density $p\in\mathcal{C}_{\geq 0}^{2}(M)$, a function $f\in\mathcal{C}^{3}(M)\cap L^{1}(M,p\,\mathrm{vol}_{g})$, a point $x\in M$, and an exponent $\eta\in(0,1/2)$.
@@ 4.2. Change of coordinates
$\Omega:=(\iota\circ\tilde{\exp}_{x})^{-1}(B_{R}(x))\subset\mathbb{R}^{d}.$
This is an open set. If $x$ is interior then $\Omega=\mathbb{B}_{R}^{d}(x)$. However, if $x$ is a border point of $M$, then $\Omega$ is a proper subset of $\mathbb{B}_{R}^{d}(x)$ admitting $0_{d}$ as a $\mathcal{C}^{0}$ boundary point.
Finally, recall that for any $y\in\mathbb{R}^{d}$ we set $y^{(k)}:=y\otimes\ldots\otimes y\in(\mathbb{R}^{d})^{\otimes^{k}}$. If $f\in\mathcal{C}^{\infty}(U)$ for some open subset $U\subset\mathbb{R}^{d}$, we let $\mathop{}\!\mathrm{d}^{(k)}_{x}f$ denote the differential of order $k$ of $f$ at $x$, which is understood here as a $k$-linear symmetric map from $(\mathbb{R}^{d})^{\otimes^{k}}$ to $\mathbb{R}$, and we set
@@
We shall also write $\mathop{}\!\mathrm{d}_{x}^{(0)}f(y^{(0)})$ for $f(x)$, in which case $\|\mathop{}\!\mathrm{d}^{(0)}_{x}f\|_{op}=f(x)$.
@@
Set $\tilde{f}:=f\circ\tilde{\exp}_{x}\circ\iota$ and $q:=(p\circ\tilde{\exp}_{x}\circ\iota)\sqrt{\det g}$, where $\sqrt{\det g}$ is the Radon–Nikodym derivative of $(\tilde{\exp}_{x}\circ\iota)^{-1}_{\#}\mathrm{vol}_{g}$ with respect to $\mathcal{L}^{d}$. Then
$\bar{L}_{t,\eta}f(x)=\frac{1}{t}\int_{\mathbb{B}^{d}_{t^{\eta}}\cap\tilde{F}_{0_{d}}(\bar{\Omega})}e^{-\frac{\|\xi\|^{2}}{t}}(\tilde{f}(0_{d})-\tilde{f}(\xi))q(\xi)\mathop{}\!\mathrm{d}\xi+o\left(\frac{1}{\sqrt{t}}\right)\qquad\text{as $t\to 0$.}$
@@
Set
[…]
$\mathbb{W}$ $:=\iota^{-1}(W).$
Then
$\bar{L}_{t,\eta}f(x)=I(t)+II(t)$
where
$I(t)$ $:=\frac{1}{t^{d/2+1}}\int_{B_{t^{\eta}}(x)\cap\tilde{\exp}_{x}(W\cap\mathbb{B}_{R}^{d})}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(x,y)}{t}\right)(f(x)-f(y))\,d\mathrm{vol}_{g}(y),$
[…]
@@
(14) $I(t)=\frac{1}{t}\int_{\mathbb{B}^{d}_{t^{\eta}}\cap\tilde{F}_{0_{d}}(\bar{\Omega})}e^{-\frac{\|\xi\|^{2}}{t}}(\tilde{f}(0_{d})-\tilde{f}(\xi))q(\xi)\mathop{}\!\mathrm{d}\xi.$
To this purpose, we point out that $W$ (and then $\mathbb{W}$) is a cone : indeed, if $v\in W$ and $\lambda>0$, then $\tilde{\gamma}_{x,\lambda v}(t)=\tilde{\gamma}_{x,v}(\lambda t)$ for any $t\in(0,t_{v}/\lambda)$, so that $\lambda v\in W$ with $t_{\lambda v}=t_{v}/\lambda$.
@@
$=\frac{1}{t^{d/2+1}}\int_{\mathbb{B}_{t^{\eta}}^{d}\cap\mathbb{W}}\exp\left(-\frac{\|\xi\|^{2}}{t}\right)(\tilde{f}(0_{d})-\tilde{f}(\xi))q(\xi)\,d\xi.$
@@
$=\frac{1}{t^{d/2+1}}\int_{(\iota\circ\tilde{\exp}_{x})^{-1}(B_{t^{\eta}}(x))\backslash\mathbb{W}}\exp\left(-\frac{\|\xi\|^{2}}{t}\right)|\tilde{f}(0)-\tilde{f}(\xi)|\tilde{q}(\xi)d\xi$
$=\frac{1}{t}\int_{\frac{(\iota\circ\tilde{\exp}_{x})^{-1}(B_{t^{\eta}}(x))}{\sqrt{t}}\backslash\mathbb{W}}\exp\left(-\|\zeta\|^{2})\right)|\tilde{f}(0_{d})-\tilde{f}(\sqrt{t}\zeta)|\tilde{q}(\sqrt{t}\zeta)d\zeta$
@@
$\tilde{q}(\sqrt{t}\zeta)=\tilde{q}(0_{d})+\mathop{}\!\mathrm{d}_{\sqrt{s_{2}}\zeta}\tilde{q}(\sqrt{t}\zeta),$
@@
$|\tilde{f}(0_{d})-\tilde{f}(\sqrt{t}\zeta)|\tilde{q}(\sqrt{t}\zeta)$ $\leq\sqrt{t}\,\tilde{q}(0_{d})|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}(\zeta)|+\frac{t\,\tilde{q}(0_{d})}{2}|\mathop{}\!\mathrm{d}^{(2)}_{\sqrt{s_{1}}\zeta}\tilde{f}(\zeta^{(2)})|$
$\phantom{=}+t\mathop{}\!\mathrm{d}_{\sqrt{s_{2}}\zeta}\tilde{q}(\zeta)|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}(\zeta)|+\frac{t^{3/2}}{2}\mathop{}\!\mathrm{d}_{\sqrt{s_{2}}\zeta}\tilde{q}(\zeta)|\mathop{}\!\mathrm{d}^{(2)}_{\sqrt{s_{1}}\zeta}\tilde{f}(\zeta^{(2)})|$
@@
$\leq\sqrt{t}\,\tilde{q}(0_{d})\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}\|_{op}\|\zeta\|+t\|\zeta\|^{2}C_{x,f,p}(t)$
@@
$C_{x,f,p}(t)$ $:=\sup_{z\in(\iota\circ\tilde{\exp}_{x})^{-1}(B_{t^{\eta}}(x))\backslash\mathbb{W}}\left(\frac{\tilde{q}(0_{d})}{2}\|\mathop{}\!\mathrm{d}^{(2)}_{z}\tilde{f}\|_{op}+\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}\|_{op}\|\mathop{}\!\mathrm{d}_{z}\tilde{q}\|_{op}\right$.
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{t^{1+\eta}}{2}\|\mathop{}\!\mathrm{d}_{z}\tilde{q}\|_{op}\|\mathop{}\!\mathrm{d}^{(2)}_{z}\tilde{f}\|_{op}\right).$
@@
$C_{x,f,p}(t)\to\frac{\tilde{q}(0_{d})}{2}\|\mathop{}\!\mathrm{d}^{(2)}_{0_{d}}\tilde{f}\|_{op}+\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}\|_{op}\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{q}\|_{op}.$
@@
Remark 4.3.
The $o(1)$ in Theorem 1.1 is likely to be improved from a deeper understanding of the asymptotic behavior of the set $\W\exp_{x}^{-1}(B_{t^{\eta}})\backslash$ W as $t\downarrow 0$.

### v3

@@ 3.3.2. Inward sector
Let us now introduce the notion of inward tangent sector for smooth manifolds with kinks. This is analogous to the inward tangent sector of Euclidean domains as discussed in Section 2.2. Our definition builds upon the classical characterization of the tangent space in terms of initial velocities of curves, which holds true for manifolds without/with boundary (see e.g. [Lee, p. 68–70]). For manifolds with boundary, one can identify inward tangent vectors at boundary points by specifying the domain of the curves we choose. This is how we came up with the following natural definition.
@@ 3.4.4. Riemannian distance and volume measure
The Riemannian distance defined via a length-minimizing problem extends with no change to the context of manifolds with kinks. We recall the definition for completeness and refer to [Lee, p.337–341], for instance, for more details.
@@ 3.5. Extension of Riemannian manifolds with kinks
For our purposes, we need to extend beyond the border any $\mathcal{C}^{2}$ Riemannian metric defined on a smooth manifold with kinks $M$. To this aim, we shall flow $M$ into its interior using a suitable semiflow. We adapt an argument for manifolds with corners that goes back to [DouadyHerault] at least, see also [Michor, Section 2.7].
@@ 3.5.2. Riemannian submanifolds with kinks
Let us now provide a definition of submanifold adapted to the context of manifolds with kinks, inspired by [Michor, p.19] who introduced submanifolds with corners.
@@
    Use Whitney’s theorem [Whitney] to extend each coordinate of $(\phi^{-1})^{*}g$ to form a $\mathcal{C}^{2}$ Riemannian metric $h$ on an open subset $V$ of $\mathbb{R}^{d}$ containing $\phi(x)$, such that $\phi^{-1}(V)\subset\tilde{U}$. Note that this extension may not be unique and depends on the choice of the border chart $(U,\phi)$.
@@ 4. Asymptotic behavior of the intrinsic Gaussian Operator
In this section, we prove Theorem 1.1 without the refined estimates on the error term $\mathrm{Err}(t)$. We refer to Section 6 for these refined estimates. Consider a smooth $d$-dimensional Riemannian manifold with kinks $M$ endowed with a $\mathcal{C}^{2}$ Riemannian metric $g$, a density $p\in\mathcal{C}_{\geq 0}^{2}(M)$, a function $f\in\mathcal{C}^{3}(M)\cap L^{1}(M,p\,\mathrm{vol}_{g})$, an exponent $\eta\in(0,1/2)$, and a point $x\in M$. For the sake of clarity, let us highlight the main steps of our proof, each of which being dedicated a subsection.
Step 1. We establish that
(12) $L_{t}f(x)=\bar{L}_{t,\eta}f(x)+O(t^{-d/2-1}e^{-t^{2\eta-1}})\qquad\text{as $t\downarrow 0$}$
with
(13) $\bar{L}_{t,\eta}f(x):=\frac{1}{t^{d/2+1}}\int_{B_{t^{\eta}}(x)}\exp\left(-\frac{\mathop{}\!\mathrm{d}_{g}^{2}(x,y)}{t}\right)(f(x)-f(y))p(y)\mathop{}\!\mathrm{d}\mathrm{vol}_{g}(y).$
Step 2. We write
(14) $\bar{L}_{t,\eta}f(x)=I(t)+II(t)$
where $I(t)$ is a term suited for an exponential change of variable, and we show that
$II(t)=o\left(\frac{1}{\sqrt{t}}\right)=:\frac{\mathrm{Err(t)}}{\sqrt{t}}\qquad\text{as $t\downarrow 0$.}$
Step 3. We prove an Euclidean version of the expansion
[…]
Step 4. We conclude by change of variable in $I(t)$.
@@ 4.1. Localisation
We perform the first step of our proof in the general context of a metric measure space $(Z,\mathop{}\!\mathrm{d},\mu)$. In this case, the intrinsic $d$-dimensional Gaussian operator at time $t>0$ associated with a density $q\in L^{1}(Z,\mu)$ is defined by
$L_{t}h(z):=\frac{1}{t^{d/2+1}}\int_{Z}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(z,y)}{t}\right)(f(z)-f(y))q(y)\mathop{}\!\mathrm{d}\mu(y)$
for any $h\in L^{1}(X,p\mu)$ and $\mu$-a.e. $z\in Z$. Then the following holds.
@@
For any $h\in L^{1}(Z,p\mu)$, $\mu$-a.e. $z\in Z$, and $t>0$,
$\left|\frac{1}{t^{d/2+1}}\int_{X\backslash B_{t^{\eta}}(z)}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(z,y)}{t}\right)(h(z)-h(y))p(y)\mathop{}\!\mathrm{d}\mathrm{\mu}(y)\right|\leq[|h(z)|\|p\|_{1}+\left\lVert hp\right\rVert_{1}]\,\frac{1}{t^{d/2+1}}e^{-t^{2\eta-1}}$
As a consequence,
$L_{t}h(z)=\bar{L}_{t}h(z)+O(t^{-d/2-1}e^{-t^{2\eta-1}})\qquad\text{as $t\downarrow 0$,}$
with
$\bar{L}_{t}h(z):=\frac{1}{t^{d/2+1}}\int_{B_{t^{\eta}}(z)}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(z,y)}{t}\right)(h(z)-h(y))p(y)\mathop{}\!\mathrm{d}\mathrm{\mu}(y).$
Proof.
By triangle inequality, for any $z,y\in X$,
$|h(z)-h(y)|\leq|h(z)||p(y)|+|hp|(y).$
Multiply by $\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(z,y)}{t}\right)$ and integrate over $y\in X\backslash B_{t^{\eta}}(z)$. The result follows from there since $\mathop{}\!\mathrm{d}^{2}(z,y)\geq t^{2\eta}$ for any such $y$. ∎
Applying the previous lemma with $(Z,\mathop{}\!\mathrm{d},\mu)=(M,\mathop{}\!\mathrm{d}_{g},\mathrm{vol}_{g})$, $h=f$, $z=x$, gives (12).
4.2. Exponential coordinates
@@
$\iota:\mathbb{R}^{d}\to T_{x}M$
denote this isometric identification. We denote by $\tilde{B}_{r}(x)$ the $\tilde{g}-$ball of radius $r$ centered at $x$, by $\tilde{\exp}_{x}$ the $\tilde{g}-$exponential map at $x$, and by $\tilde{\gamma}_{x,v}$ the unique maximal $\tilde{g}-$geodesic with initial point $x$ and initial velocity $v\in T_{x}\tilde{M}$.
Since $x$ is an interior point for $\tilde{M}$, there exists $R>0$ such that $\tilde{\exp}_{x}$ is a diffeomorphism from $\{\|\cdot\|_{\tilde{g}_{x}}<R\}\subset T_{x}\tilde{M}$ onto $\tilde{B}_{R}(x)$. This implies, in particular, that $(B_{R}(x),(\iota\circ\tilde{\exp}_{x})^{-1})$ is a $d$-dimensional chart centered at $x$ of $M$. This chart is interior (resp. border) if $x$ is interior (resp. border) for $M$.
@@
Consider the open subset of $\mathbb{R}^{d}$ defined as
(16) $\Omega:=(\iota\circ\tilde{\exp}_{x})^{-1}(B_{R}(x)).$
Note that if $x$ is interior then $\Omega=\mathbb{B}_{R}^{d}$. However, if $x$ is border, then $\Omega$ is a proper subset of $\mathbb{B}_{R}^{d}$ admitting $0_{d}$ as a $\mathcal{C}^{0}$ boundary point. Moreover, the regularity of $x$ as a border point transfers to the regularity of $0_{d}$ as a $\mathcal{C}^{0}$ boundary point of $\Omega$.
@@
$\mathbb{W}$ $:=\iota^{-1}(W),$
and for any $t>0$ we introduce
$I(t)$ $:=\frac{1}{t^{d/2+1}}\int_{B_{t^{\eta}}(x)\cap\tilde{\exp}_{x}(W\cap\mathbb{B}_{R}^{d})}\exp\left(-\frac{\mathop{}\!\mathrm{d}^{2}(x,y)}{t}\right)(f(x)-f(y))\,p(y)\,d\mathrm{vol}_{g}(y),$
[…]
Obviously,
$\bar{L}_{t}f(x)=I(t)+II(t).$
Moreover, $W$ is a cone : indeed, if $v\in W$ and $\lambda>0$, then $\tilde{\gamma}_{x,\lambda v}(t)=\tilde{\gamma}_{x,v}(\lambda t)$ for any $t\in(0,t_{v}/\lambda)$, so that $\lambda v\in W$ with $t_{\lambda v}=t_{v}/\lambda$. This obviously implies that $\mathbb{W}$ is a cone too.
Finally, recall that for any $y\in\mathbb{R}^{d}$ we set $y^{(k)}:=y\otimes\ldots\otimes y\in(\mathbb{R}^{d})^{\otimes^{k}}$. If $h\in\mathcal{C}^{\infty}(U)$ for some open subset $U\subset\mathbb{R}^{d}$, we let $\mathop{}\!\mathrm{d}^{(k)}_{z}h$ denote the differential of order $k$ of $h$ at $z\in\mathbb{R}^{d}$, which is understood here as a $k$-linear symmetric map from $(\mathbb{R}^{d})^{\otimes^{k}}$ to $\mathbb{R}$, and we set
[…]
We shall also write $\mathop{}\!\mathrm{d}_{z}^{(0)}f(y^{(0)})$ for $f(z)$, in which case $\|\mathop{}\!\mathrm{d}^{(0)}_{z}f\|_{op}=f(z)$.
@@
Set $\tilde{f}:=f\circ\tilde{\exp}_{x}\circ\iota$ and $\tilde{p}:=(p\circ\tilde{\exp}_{x}\circ\iota)\sqrt{\det g}$, where $\sqrt{\det g}$ is the Radon–Nikodym derivative of $(\tilde{\exp}_{x}\circ\iota)^{-1}_{\#}\mathrm{vol}_{g}$ with respect to $\mathcal{L}^{d}$. Then
$\bar{L}_{t,\eta}f(x)=\frac{1}{t}\int_{\mathbb{B}^{d}_{t^{\eta}}\cap\tilde{F}_{0_{d}}(\bar{\Omega})}e^{-\frac{\|\xi\|^{2}}{t}}(\tilde{f}(0_{d})-\tilde{f}(\xi))\tilde{p}(\xi)\mathop{}\!\mathrm{d}\xi+o\left(\frac{1}{\sqrt{t}}\right)\qquad\text{as $t\to 0$.}$
@@
(20) $I(t)=\frac{1}{t}\int_{\mathbb{B}^{d}_{t^{\eta}}\cap\tilde{F}_{0_{d}}(\bar{\Omega})}e^{-\frac{\|\xi\|^{2}}{t}}(\tilde{f}(0_{d})-\tilde{f}(\xi))\tilde{p}(\xi)\mathop{}\!\mathrm{d}\xi.$
@@
$=\frac{1}{t^{d/2+1}}\int_{\mathbb{B}_{t^{\eta}}^{d}\cap\mathbb{W}}\exp\left(-\frac{\|\xi\|^{2}}{t}\right)(\tilde{f}(0_{d})-\tilde{f}(\xi))\tilde{p}(\xi)\,d\xi.$
@@
(26) $=\frac{1}{t^{d/2+1}}\int_{(\iota\circ\tilde{\exp}_{x})^{-1}(B_{t^{\eta}}(x))\backslash\mathbb{W}}\exp\left(-\frac{\|\xi\|^{2}}{t}\right)|\tilde{f}(0)-\tilde{f}(\xi)|\tilde{p}(\xi)d\xi$
$=\frac{1}{t}\int_{\frac{(\iota\circ\tilde{\exp}_{x})^{-1}(B_{t^{\eta}}(x))}{\sqrt{t}}\backslash\mathbb{W}}\exp\left(-\|\zeta\|^{2})\right)|\tilde{f}(0_{d})-\tilde{f}(\sqrt{t}\zeta)|\tilde{p}(\sqrt{t}\zeta)d\zeta$
@@
$\tilde{p}(\sqrt{t}\zeta)=\tilde{p}(0_{d})+\mathop{}\!\mathrm{d}_{\sqrt{s_{2}}\zeta}\tilde{p}(\sqrt{t}\zeta),$
@@
$|\tilde{f}(0_{d})-\tilde{f}(\sqrt{t}\zeta)|\tilde{p}(\sqrt{t}\zeta)$ $\leq\sqrt{t}\,\tilde{p}(0_{d})|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}(\zeta)|+\frac{t\,\tilde{p}(0_{d})}{2}|\mathop{}\!\mathrm{d}^{(2)}_{\sqrt{s_{1}}\zeta}\tilde{f}(\zeta^{(2)})|$
$\phantom{=}+t\mathop{}\!\mathrm{d}_{\sqrt{s_{2}}\zeta}\tilde{p}(\zeta)|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}(\zeta)|+\frac{t^{3/2}}{2}\mathop{}\!\mathrm{d}_{\sqrt{s_{2}}\zeta}\tilde{p}(\zeta)|\mathop{}\!\mathrm{d}^{(2)}_{\sqrt{s_{1}}\zeta}\tilde{f}(\zeta^{(2)})|$
@@
$\leq\sqrt{t}\,\tilde{p}(0_{d})\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}\|_{op}\|\zeta\|+t\|\zeta\|^{2}C_{x,f,p}(t)$
@@
$C_{x,f,p}(t)$ $:=\sup_{z\in(\iota\circ\tilde{\exp}_{x})^{-1}(B_{t^{\eta}}(x))\backslash\mathbb{W}}\left(\frac{\tilde{p}(0_{d})}{2}\|\mathop{}\!\mathrm{d}^{(2)}_{z}\tilde{f}\|_{op}+\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}\|_{op}\|\mathop{}\!\mathrm{d}_{z}\tilde{p}\|_{op}\right$.
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\left.+\frac{t^{1+\eta}}{2}\|\mathop{}\!\mathrm{d}_{z}\tilde{p}\|_{op}\|\mathop{}\!\mathrm{d}^{(2)}_{z}\tilde{f}\|_{op}\right).$
@@
$C_{x,f,p}(t)\to\frac{\tilde{p}(0_{d})}{2}\|\mathop{}\!\mathrm{d}^{(2)}_{0_{d}}\tilde{f}\|_{op}+\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{f}\|_{op}\|\mathop{}\!\mathrm{d}_{0_{d}}\tilde{p}\|_{op}.$